OpenStudy (anonymous):

Two trains leave stations 448 miles apart at the same time and travel toward each other. One train travels at 85 miles per hour while the other travels at 75 miles per hour. How long will it take for the two trains to meet? Do not do any rounding.

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{time}=\dfrac{\normalsize \text{distance}}{\normalsize \text{relative speed}}\hspace{.33em}\\~\\ \normalsize \text{relative speed}=(s_1+s_2)\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

am getting 5.27 and 5.97

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{time}=\dfrac{448}{85+75}=2.8\quad hrs\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

ohhh i get it thank you sir